OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. (−6 + i) − (5 − 2i)

OpenStudy (anonymous):

Just subtract them. You have -6 ones and you subtract 5 ones. How many do you have left? You have 1 i and you subtract -2 i. How many do you have left?

OpenStudy (anonymous):

are u saying its 11 and 3i

OpenStudy (anonymous):

-6 - 5 = -11 1i - (-2)i = 3i So (-6 + i) - (5-2i) = (-6 - 5) + (1i - (-2)i) = -11 + 3i

OpenStudy (anonymous):

These really aren't difficult, you just have to make sure you understand what is being asked of you.